OpenStudy (shealah):

what is an equation of the line that passes through the point (7,3) and is parallel to the line 4x+2y=10?

OpenStudy (shealah):

I put that into y=mx+b form and got y= -2x+5

jimthompson5910 (jim_thompson5910):

what's the slope of y= -2x+5

OpenStudy (shealah):

-2

jimthompson5910 (jim_thompson5910):

So this unknown parallel line will also have a slope of -2

jimthompson5910 (jim_thompson5910):

y = mx+b y = -2x+b Now plug in the given point (x,y) = (7,3) 3 = -2(7) + b then solve for b

OpenStudy (shealah):

oh okay. Thanks :)

jimthompson5910 (jim_thompson5910):

yw